# MedFLIP: Medical Vision-and-Language Self-supervised Fast Pre-Training with Masked Autoencoder

### v1

@@ Abstract
Within the domain of medical analysis, extensive research has explored the potential of mutual learning between Masked Autoencoders(MAEs) and multimodal data. However, the impact of MAEs on intermodality remains a key challenge. We introduce MedFLIP, a Fast Language-Image Pre-training method for Medical analysis. We explore MAEs for zero-shot learning with crossed domains, which enhances the model’s ability to learn from limited data, a common scenario in medical diagnostics. We verify that masking an image does not affect inter-modal learning. Furthermore, we propose the SVD loss to enhance the representation learning for characteristics of medical images, aiming to improve classification accuracy by leveraging the structural intricacies of such data. Lastly, we validate using language will improve the zero-shot performance for the medical image analysis. MedFLIP’s scaling of the masking process marks an advancement in the field, offering a pathway to rapid and precise medical image analysis without the traditional computational bottlenecks. Through experiments and validation, MedFLIP demonstrates efficient performance improvements, setting an explored standard for future research and application in medical diagnostics.
@@ 1 Introduction
Figure 1: Our proposed method, MedFLIP, demonstrates a superior trade-off between training efficiency and accuracy compared to the MedCLIP method. Notably, MedFLIP achieves higher accuracy on the CheXpert-5x200 [7] validation set with zero-shift evaluation, while maintaining a consistent model size.
@@
The annotation of medical data presents huge challenges, necessitating considerable effort to enhance its accuracy and efficiency. Due to the adaptability of MAE [5] to natural images, many works have used the idea of MAEs for medical images. In light of this, there has been a burgeoning interest in the exploration of self-supervised learning approaches for medical image data. Given the inherent structural similarities across medical images, a substantial body of research has focused on the development and application of image masking techniques[9, 22, 28] within this domain. This focus underscores the potential of self-supervised learning to leverage the unique characteristics of medical images, facilitating advancements in the automated analysis and interpretation of such data.
@@
To simplify the process of training models for medical images, leveraging natural language supervision offers a promising route by enabling a richer contextual understanding of images through the relationships between objects, scenes, actions, and contexts. The principle of masking, as seen in models like FLIP [14], can be adapted to efficiently handle vast amounts of medical imaging data. By applying masking techniques, we can process a larger number of sample pairs within the same time frame and with similar memory requirements. This approach allows for more efficient comparison and contrast of sample pairs, enhancing the training’s effectiveness without significantly increasing computational demands. Specifically for medical images, adopting such methods could significantly reduce training time and improve accuracy, providing a balanced trade-off that benefits both energy consumption and the quality of the model’s predictions. This streamlined training approach could unlock new potentials in medical image analysis, making it a valuable area for further research and application.
@@
In striving to minimize dependency on extensive datasets and reduce training durations, advancements in vision and language integration, exemplified by models like CLIP [18], suggest that textual information significantly enhances image comprehension and image-text retrieval. This synergy between textual and visual data has been instrumental in facilitating more efficient model training processes. Consequently, this multimodal approach is being explored within the realm of medical imaging [15, 21, 23], proposing an innovative methodology that leverages textual annotations to augment the understanding and analysis of medical images. Such endeavors aim to harness the complementary strengths of vision and language, potentially revolutionizing the efficiency and efficacy of medical image analysis by capitalizing on the inherent relationship between text and image for improved model performance with less reliance on large datasets.
@@
However, the application of AI [18] in this domain faces unique challenges, primarily due to the complexity and variability of medical data. Traditional machine learning models require large datasets for training to achieve high accuracy, which is often a significant barrier in the medical field where data can be scarce, sensitive, and expensive to annotate. This scarcity of large, annotated datasets has necessitated the development of models capable of learning effectively from limited data. In this context, our work introduces MedFLIP, a fast vision-language pipeline designed for medical image analysis by addressing these critical challenges through innovative few-shot learning capabilities and a novel loss function.
@@ 2 Related Work
MAEs are a type of self-supervised learning method for vision, introduced by Kaiming He et al. [5], which have shown remarkable effectiveness in learning representations from unlabelled data by predicting masked parts of the input images [3, 9, 22, 28, 1, 4]. Building upon this foundational work, MAEs have been adapted for medical imaging, demonstrating significant potential in enhancing diagnostic accuracy and efficiency. In the medical domain, MAEs leverage the inherent structure and patterns within medical images, such as MRI and CT scans, to learn robust features that can aid in disease detection, segmentation, and even prognosis prediction. This adaptation of MAEs to medical imaging not only underscores the versatility of self-supervised learning methods but also opens new avenues for research and application in healthcare, potentially transforming the way medical images are analyzed and interpreted.
@@
The evolution of multimodal learning has seen the emergence of models that can understand and relate information across different data types. CLIP (Contrastive Language-Image Pretraining) [18] has set a precedent by efficiently learning visual concepts from natural language supervision. Some Multimodel work [11, 12, 24, 1] delve into multimodal applications. Models pre-trained on ImageNet have been repurposed with fine-tuning to achieve impressive results in tasks such as X-ray and MRI interpretation [19]. In the medical domain, efforts like ConVIRT [26] extend this approach to bridge the gap between visual and textual medical data, enhancing the understanding of radiology images in conjunction with associated reports [26]. GLoRIA [6] and MedCLIP [21] follow in these footsteps, with MedCLIP, in particular, pushing the boundaries by directly training on raw medical image-text pairs and achieving state-of-the-art results in image-text retrieval tasks on datasets such as CheXpert5x200 [8, 7].
@@
Zero-shot learning in imaging [16, 20, 2, 13, 27, 10] presents a promising avenue for addressing the challenge of limited labeled data by leveraging auxiliary information from related domains. Recent advancements in ZSL techniques have shown potential in transferring knowledge across domains for tasks such as disease diagnosis, organ segmentation, and anomaly detection. For instance, approaches integrating generative adversarial networks (GANs) with ZSL frameworks have demonstrated robustness in learning representations from unannotated medical images by aligning feature distributions between seen and unseen classes. Furthermore, methods incorporating semantic embeddings and knowledge graphs have facilitated knowledge transfer between domains, enhancing the generalization capability of models in medical image analysis tasks  [16, 17, 15].
@@ 3.2 Masked Autoencoder
Our methods are inspired by MAE [5]. For medical image analysis, we utilize transformer-based methods from MedClip [21] and MAE [5]. We observe that MAE’s transformer architecture maintains the same token number across different mask ratios, ensuring consistency in our approach.
@@ 4 Experiments
Figure 3: In the context of zero-shot learning, our proposed method, MedFLIP, exhibits superior performance compared to established approaches, namely MedCLIP, ConVIRT [25], and GLoRIA [6]. This is particularly evident when utilizing limited pretraining data. While ConVIRT and GLoRIA leverage the MIMIC-CXR (369K) and CheXpert (191K) datasets, respectively, MedFLIP demonstrates demonstrably improved performance.

### v2

@@ Abstract
Within the domain of medical analysis, extensive research has explored the potential of mutual learning between Masked Autoencoders(MAEs) and multimodal data. However, the impact of MAEs on intermodality remains a key challenge. We introduce MedFLIP, a Fast Language-Image Pre-training method for Medical analysis. We explore MAEs for zero-shot learning with crossed domains, which enhances the model’s ability to learn from limited data, a common scenario in medical diagnostics. We verify that masking an image does not affect inter-modal learning. Furthermore, we propose the SVD loss to enhance the representation learning for characteristics of medical images, aiming to improve classification accuracy by leveraging the structural intricacies of such data. Our theory posits that masking encourages semantic preservation, robust feature extraction, regularization, domain adaptation, and invariance learning. Lastly, we validate using language will improve the zero-shot performance for the medical image analysis. MedFLIP’s scaling of the masking process marks an advancement in the field, offering a pathway to rapid and precise medical image analysis without the traditional computational bottlenecks. Through experiments and validation, MedFLIP demonstrates efficient performance improvements, helps for future research and application in medical diagnostics.
@@ 1 Introduction
Figure 1: Our proposed method, MedFLIP, demonstrates a superior trade-off between training efficiency and accuracy compared to the MedCLIP method. Notably, MedFLIP achieves higher accuracy on the CheXpert-5x200 [irvin2019chexpert] validation set with zero-shift evaluation, while maintaining a consistent model size.
@@
The annotation of medical data presents huge challenges, necessitating considerable effort to enhance its accuracy and efficiency. Due to the adaptability of MAE [he2022masked] to natural images, many works have used the idea of MAEs for medical images. In light of this, there has been a burgeoning interest in the exploration of self-supervised learning approaches for medical image data. Given the inherent structural similarities across medical images, a substantial body of research has focused on the development and application of image masking techniques[ke2022mask, xiao2023delving, zhou2023self] within this domain. This focus underscores the potential of self-supervised learning to leverage the unique characteristics of medical images, facilitating advancements in the automated analysis and interpretation of such data.
@@
To simplify the process of training models for medical images, leveraging natural language supervision offers a promising route by enabling a richer contextual understanding of images through the relationships between objects, scenes, actions, and contexts. The principle of masking, as seen in models like FLIP [li2023scaling], can be adapted to efficiently handle vast amounts of medical imaging data. By applying masking techniques, we can process a larger number of sample pairs within the same time frame and with similar memory requirements. This approach allows for more efficient comparison and contrast of sample pairs, enhancing the training’s effectiveness without significantly increasing computational demands. Specifically for medical images, adopting such methods could significantly reduce training time and improve accuracy, providing a balanced trade-off that benefits both energy consumption and the quality of the model’s predictions. This streamlined training approach could unlock new potentials in medical image analysis, making it a valuable area for further research and application.
@@
In striving to minimize dependency on extensive datasets and reduce training durations, advancements in vision and language integration, exemplified by models like CLIP [radford2021learning], suggest that textual information significantly enhances image comprehension and image-text retrieval. This synergy between textual and visual data has been instrumental in facilitating more efficient model training processes. Consequently, this multimodal approach is being explored within the realm of medical imaging [liu2023clip, wang2022medclip, ye2023uniseg], proposing an innovative methodology that leverages textual annotations to augment the understanding and analysis of medical images. Such endeavors aim to harness the complementary strengths of vision and language, potentially revolutionizing the efficiency and efficacy of medical image analysis by capitalizing on the inherent relationship between text and image for improved model performance with less reliance on large datasets.
@@
However, the application of AI [radford2021learning] in this domain faces unique challenges, primarily due to the complexity and variability of medical data. Traditional machine learning models require large datasets for training to achieve high accuracy, which is often a significant barrier in the medical field where data can be scarce, sensitive, and expensive to annotate. This scarcity of large, annotated datasets has necessitated the development of models capable of learning effectively from limited data. In this context, our work introduces MedFLIP, a fast vision-language pipeline designed for medical image analysis by addressing these critical challenges through innovative few-shot learning capabilities and a novel loss function.
@@ 2 Related Work
MAEs are a type of self-supervised learning method for vision, introduced by Kaiming He et al. [he2022masked], which have shown remarkable effectiveness in learning representations from unlabelled data by predicting masked parts of the input images [feichtenhofer2022masked, ke2022mask, xiao2023delving, zhou2023self, bachmann2022multimae, gupta2024siamese]. Building upon this foundational work, MAEs have been adapted for medical imaging, demonstrating significant potential in enhancing diagnostic accuracy and efficiency. In the medical domain, MAEs leverage the inherent structure and patterns within medical images, such as MRI and CT scans, to learn robust features that can aid in disease detection, segmentation, and even prognosis prediction. This adaptation of MAEs to medical imaging not only underscores the versatility of self-supervised learning methods but also opens new avenues for research and application in healthcare, potentially transforming the way medical images are analyzed and interpreted.
@@
The evolution of multimodal learning has seen the emergence of models that can understand and relate information across different data types. CLIP (Contrastive Language-Image Pretraining) [radford2021learning] has set a precedent by efficiently learning visual concepts from natural language supervision. Some Multimodel work [li2024cpseg, li2024tree, yu2022metaformer, bachmann2022multimae] delve into multimodal applications. Models pre-trained on ImageNet have been repurposed with fine-tuning to achieve impressive results in tasks such as X-ray and MRI interpretation [rajpurkar2017chexnet]. In the medical domain, efforts like ConVIRT [zhang2020convirt] extend this approach to bridge the gap between visual and textual medical data, enhancing the understanding of radiology images in conjunction with associated reports [zhang2020convirt]. GLoRIA [huang2021gloria] and MedCLIP [wang2022medclip] follow in these footsteps, with MedCLIP, in particular, pushing the boundaries by directly training on raw medical image-text pairs and achieving state-of-the-art results in image-text retrieval tasks on datasets such as CheXpert5x200 [johnson2019mimic, irvin2019chexpert].
@@
Zero-shot learning in imaging [mahapatra2021medical, rezaei2020zero, chen2023cunerf, li2023mask, zhou2023multi, li2023segment] presents a promising avenue for addressing the challenge of limited labeled data by leveraging auxiliary information from related domains. Recent advancements in ZSL techniques have shown potential in transferring knowledge across domains for tasks such as disease diagnosis, organ segmentation, and anomaly detection. For instance, approaches integrating generative adversarial networks (GANs) with ZSL frameworks have demonstrated robustness in learning representations from unannotated medical images by aligning feature distributions between seen and unseen classes. Furthermore, methods incorporating semantic embeddings and knowledge graphs have facilitated knowledge transfer between domains, enhancing the generalization capability of models in medical image analysis tasks  [mahapatra2021medical, mumuni2023improving, liu2023clip].
@@ 3.2 Masked Autoencoder
Our methods are inspired by MAE [he2022masked]. For medical image analysis, we utilize transformer-based methods from MedClip [wang2022medclip] and MAE [he2022masked]. We observe that MAE’s transformer architecture maintains the same token number across different mask ratios, ensuring consistency in our approach.
@@ 3.3 Fusion Module
3.4 Theoretical Analysis
Assumption 1.
The masked autoencoder in MedFLIP is trained on a dataset of medical image-text pairs $(x_{i},t_{i})$ drawn i.i.d. from an unknown joint distribution $\mathcal{D}$.
Assumption 2.
The SVD loss $L_{SVD}$ is $L$-Lipschitz continuous with respect to the singular value $\sigma_{1}$ of the similarity matrix $S$ between image and text embeddings.
Lemma 1.
Let $\hat{y}_{ij}$ be the predicted normalized similarity between the $i$-th image embedding and $j$-th text embedding in a batch (Eq. 5 in the paper). Let $y^{*}_{ij}$ be the optimal normalized similarity that minimizes the MedFLIP loss $L_{MedFLIP}$ (Eq. 7). Then with probability at least $1-\delta$, we have:
$|\hat{y}_{ij}-y^{*}_{ij}|\leq\frac{1}{\tau}\sqrt{\frac{\log(2/\delta)}{2N_{batch}}}+\frac{\beta L}{\tau}$
where $\tau$ is the temperature hyperparameter and $\beta$ is the SVD loss weight.
Proof.
The first term follows from applying Hoeffding’s inequality to bound the deviation between the empirical mean $\hat{y}_{ij}$ and true expectation $y^{*}_{ij}$ of the normalized similarity, using the fact that $0\leq\hat{y}_{ij}\leq 1$. The second term follows from the Lipschitz continuity of $L_{SVD}$. Specifically, by Assumption 2, for any two singular values $\sigma_{1},\sigma_{1}^{\prime}$, we have $|L_{SVD}(\sigma_{1})-L_{SVD}(\sigma_{1}^{\prime})|\leq L|\sigma_{1}-\sigma_{1}^{\prime}|$. The deviation $|\sigma_{1}-\sigma_{1}^{|}$ between the singular value for $\hat{y}$ vs $y^{*}$ is bounded by their total variation distance, which is at most twice the bound on $|\hat{y}_{ij}-y^{*}_{ij}|$ from the first term. Combining this with the Lipschitz property and the $\beta/\tau$ factor completes the proof. ∎
Theorem 3.1 (MedFLIP Generalization Bound).
Let $f_{\theta}$ denote the MedFLIP model with learned parameters $\theta$. Let $\hat{R}(f_{\theta})$ and $R(f_{\theta})$ denote its empirical and expected risks respectively on a downstream task. Then with probability at least $1-\delta$ over the training set, we have:
$R(f_{\theta})\leq\hat{R}(f_{\theta})+\mathcal{O}\left(\frac{1}{\tau}\sqrt{\frac{\log(1/\delta)}{N_{batch}}}+\frac{\beta L}{\tau}\right)$
Proof.
The empirical risk $\hat{R}(f_{\theta})$ is an average over the normalized similarity terms $\hat{y}_{ij}$ for aligned image-text pairs. By Lemma 1 and a union bound over all $O(N{batch}^{2})$ pairs, each term $\hat{y}_{ij}$ concentrates around the optimal $y^{*}_{ij}$ with probability at least $1-\delta$, up to an error of $\frac{1}{\tau}\sqrt{\frac{\log(2N_{batch}^{2}/\delta)}{2N_{batch}}}+\frac{\beta L}{\tau}$. When this holds, the deviation $|\hat{R}(f_{\theta})-R(f_{\theta})|$ is also bounded by the same error term. Simplifying the logarithmic factor and suppressing constants gives the stated bound. ∎
The theorem bounds the generalization error of the MedFLIP model on a downstream task, in terms of the number of image-text pairs per training batch, the temperature parameter $\tau$, the SVD loss weight $\beta$, and the Lipschitz constant $L$ of the SVD loss. The key assumptions are that the training data is i.i.d. and that the SVD loss is Lipschitz continuous.
@@ 4 Experiments
Figure 3: In the context of zero-shot learning, our proposed method, MedFLIP, exhibits superior performance compared to established approaches, namely MedCLIP, ConVIRT [zhang2022contrastive], and GLoRIA [huang2021gloria]. This is particularly evident when utilizing limited pretraining data. While ConVIRT and GLoRIA leverage the MIMIC-CXR (369K) and CheXpert (191K) datasets, respectively, MedFLIP demonstrates demonstrably improved performance.